# FLUKE: A Linguistically-Driven and Task-Agnostic Framework for Robustness Evaluation

### v2

@@ A.3 Examples of annotation decisions for generative tasks (simplified quality assurance scenario)
We provide examples of decisions where we kept the modification together with the original label for the task, changed the label, or rejected the modification, in Table˜8.
@@ A.4 LLM settings
We report the exact endpoint and sampling parameters for LLMs in Table˜9.
@@ A.5 PLM settings
We report the finetuning details for PLMs in Table˜10.
@@
Fine-tuning details are in Table˜10.
@@
A.6 Use Case: Dialogue Contradiction Detection and Sentiment Analysis
As Table˜11 shows, on Dialogue contradiction detection task base LLMs are overall more robust to modifications than PLMs; however, reasoning LLMs on average perform similar to LLMs. All models show major brittleness to negation, and the majority of them also lack robustness to change of grammatical roles, which shows that they tend to rely on surface-level, token consistency between the context and the final turn, rather than on logical coherence. Moreover, LLMs (but to a lesser extent PLMs) demonstrate geographical bias, whuile PLMs are particularly brittle to casual style.
Category Modification PLM LLM Avg
BERT GPT2 T5 GPT4o Claude Llama GPT5 DS Avg
Bias Temporal 8.0 6.0 7.0 2.0 5.0 1.0 3.0 2.0 4.2
Geographical 6.5 8.7 12.0 13.0 15.2 12.0 18.5 13.0 12.4
Length 10.0 6.0 5.0 4.0 6.0 1.0 6.0 7.0 5.6
Orthogr. Spelling 6.0 6.0 4.0 8.0 3.0 2.0 1.0 6.0 4.5
Capitalization 0.0 6.2 6.2 3.1 2.1 2.1 4.2 1.0 3.1
Punctuation 4.0 2.0 8.0 3.0 3.0 1.0 5.0 2.0 3.5
Morpho. Derivation 0.0 2.2 1.1 4.3 5.4 1.1 5.4 6.5 3.2
Compound 6.0 4.0 3.0 4.0 7.0 1.0 8.0 6.0 4.9
Syntax Voice 9.0 11.0 7.0 7.0 2.0 3.0 9.0 5.0 6.6
Grammar 14.3 12.9 7.1 20.6 7.4 11.8 16.2 19.1 13.7
Conjunction 5.0 4.0 7.0 6.0 1.0 2.0 4.0 4.0 4.1
Semantics Concept 7.0 5.0 4.0 8.0 7.0 2.0 10.0 6.0 6.1
Negation 29.0 39.0 31.0 31.0 30.0 30.0 25.0 25.0 30.0
Discourse Disc. markers 5.7 9.2 3.4 3.4 2.3 3.4 5.7 5.7 4.9
Appraisal 6.0 10.0 5.0 7.0 6.0 4.0 10.0 11.0 7.4
Varieties Style 12.0 10.0 12.0 6.0 4.0 2.0 6.0 6.0 7.2
Dialect 20.8 5.7 9.4 11.3 15.1 7.5 5.7 13.2 11.1
Average 8.8 8.7 7.8 8.3 7.1 5.1 8.4 8.2 7.8
Table 11: Dialogue: Unrobustness (U, %) by model and modification. Claude stands for Claude 3.5, Llama for Llama 3.1, and DS for DeepSeek R1
On Sentiment Analysis task, overall, PLMs were less robust to modifications compared to LLMs (see Table˜12), with the exception of Claude-3.5 which lacks robusteness to many tests that the other LLMs are stable on. Most prominently, all LLMs and PLMs are significantly unrobust to negation, and can be easily misguided by adding another appraisal marker to a sentence already bearing some sentiment.
Category Modification PLM LLM
BERT GPT2 T5 GPT4o Claude Llama GPT5 DS Avg
Bias Temporal 5.0 5.0 3.0 0.0 8.0 1.0 3.0 3.0 3.5
Geographical 5.0 4.0 3.0 9.0 8.0 7.0 4.0 8.0 6.0
Length 3.0 2.0 3.0 3.0 9.0 2.0 1.0 1.0 3.0
Orthogr. Spelling 4.0 4.0 0.0 0.0 6.0 3.0 0.0 0.0 2.1
Capitalization 0.0 3.0 3.0 1.0 7.1 3.0 0.0 3.0 2.5
Punctuation 1.0 2.0 0.0 1.0 3.0 3.0 1.0 1.0 1.5
Morphol. Derivation 3.4 5.7 2.3 2.3 6.9 2.3 3.4 4.6 3.9
Compound 3.2 8.4 3.2 2.1 5.3 1.1 3.2 4.2 3.8
Syntactic Voice 9.0 10.0 9.0 4.0 10.0 3.0 5.0 4.0 6.8
Grammar 4.5 12.1 3.0 4.5 7.6 4.5 4.5 4.5 5.7
Conjunction 3.0 6.0 1.0 1.0 3.0 3.0 0.0 0.0 2.1
Semantic Concept 6.0 6.0 5.0 4.0 4.0 2.0 1.0 5.0 4.1
Negation 22.9 20.8 25.0 16.7 17.7 15.6 16.7 17.7 19.1
Pragmatic Discourse 3.0 6.1 4.0 1.0 12.1 3.0 2.0 3.0 4.3
Sentiment 19.0 16.0 12.0 14.0 15.0 16.0 11.0 10.0 14.1
Genre Casual 9.0 9.0 5.0 3.0 7.0 6.0 3.0 2.0 5.5
Dialectal 9.8 9.8 7.8 4.9 6.9 5.9 2.9 3.9 6.5
Average 6.5 7.6 5.3 4.2 8.0 4.8 3.6 4.4 5.6
Table 12: Sentiment Analysis: Unrobustness (U, %) by model and modification. Claude stands for Claude 3.5, Llama for Llama 3.1, and DS for DeepSeek R1
A.7 Explicitly making model aware of the modification
@@
A.8 Unrobustness results for each model averaged across tasks
Table˜14 reports the Unrobustness scores for each model averaged across all tasks. Note that IFEval and GSM tasks only contain LLMs results. Overall, PLMs are more unrobust compared to LLMs. Among LLMs, the newer reasoning models are more robust compared to the base versions. Modification-wise, Geographical Bias, Negation, Style, and Dialect are the most prominent ones in creating instability.

### v3

@@ 8 Limitations
Acknowledgements
This research is supported by Oracle. Lau was supported by Australian Research Council
@@ A.3 Examples of annotation decisions for generative tasks (simplified quality assurance scenario)
We provide examples of decisions where we kept the modification together with the original label for the task, changed the label, or rejected the modification, in Table˜9.
@@ A.4 LLM settings
We report the exact endpoint and sampling parameters for LLMs in Table˜10.
@@ A.5 PLM settings
We report the finetuning details for PLMs in Table˜11.
@@
Fine-tuning details are in Table˜11.
@@
A.6 Explicitly making model aware of the modification
@@
A.7 Use Case: NER and Coreference tasks (extended results)
We present the full results of NER and Coreference tasks with confidence interval in Table˜13 and Table˜14.
Category Modification PLM LLM
BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 3.7 [1.4, 6.9] 3.0 [0.4, 6.4] 1.2 [0.3, 2.6] 6.2 [2.5, 10.8] 4.3 [1.1, 8.2] 1.8 [0.3, 4.0] 7.7 [4.4, 11.5] 10.6 [6.4, 15.4] 4.8 [2.1, 8.2]
Geographical 25.1 [19.6, 30.9] 27.6 [21.7, 33.9] 29.0 [23.6, 34.5] 22.3 [17.5, 27.5] 26.0 [19.5, 32.8] 27.0 [21.5, 32.6] 22.0 [16.7, 27.7] 27.8 [22.7, 33.0] 25.8 [20.4, 31.6]
Length 11.8 [6.6, 17.7] 12.2 [7.4, 17.7] 13.1 [8.5, 18.3] 12.7 [7.6, 18.5] 15.7 [9.3, 23.1] 6.4 [3.0, 10.5] 9.7 [5.2, 15.1] 12.6 [7.8, 18.0] 11.8 [6.9, 17.4]
Orthography Spelling 4.3 [1.6, 7.7] 2.5 [0.4, 5.2] 1.4 [0.5, 2.5] 6.6 [3.8, 9.8] 3.0 [0.5, 6.2] 3.9 [1.6, 6.6] 7.3 [4.3, 10.8] 11.4 [7.3, 15.9] 5.0 [2.5, 8.1]
Capitalization 0.9 [0.1, 1.8] 19.3 [13.1, 26.1] 13.1 [8.5, 18.2] 11.3 [6.6, 16.6] 8.9 [4.5, 14.1] 15.3 [9.7, 21.6] 9.1 [5.6, 13.1] 13.6 [9.1, 18.5] 11.5 [7.2, 16.3]
Punctuation 6.5 [3.0, 10.7] 3.7 [1.3, 6.8] 4.9 [2.4, 7.8] 7.1 [3.6, 11.2] 9.1 [5.0, 13.7] 7.9 [4.1, 12.4] 7.8 [4.5, 11.7] 15.0 [10.3, 20.2] 7.8 [4.3, 11.8]
Morphology Derivation 1.9 [0.5, 3.7] 4.2 [1.4, 7.5] 5.8 [2.1, 10.5] 3.7 [1.1, 7.2] 2.0 [0.1, 4.5] 1.3 [0.1, 3.2] 9.5 [4.4, 15.7] 8.2 [4.2, 12.9] 4.6 [1.7, 8.1]
Compound 3.1 [0.8, 6.1] 0.6 [0.0, 1.7] 5.2 [2.0, 9.1] 3.1 [0.8, 6.5] 1.7 [0.0, 4.4] 1.0 [0.0, 2.7] 7.0 [3.4, 11.2] 12.9 [8.2, 18.1] 4.3 [1.9, 7.5]
Syntax Voice 7.8 [3.9, 12.4] 10.8 [6.0, 16.2] 5.7 [3.1, 8.7] 7.5 [3.6, 12.2] 5.5 [2.1, 9.7] 4.5 [1.5, 8.2] 8.3 [4.3, 13.1] 11.3 [7.2, 16.0] 7.7 [4.0, 12.1]
Grammar 8.0 [4.0, 12.9] 15.9 [10.4, 21.8] 10.5 [6.3, 15.5] 8.3 [4.7, 12.5] 3.5 [1.2, 6.6] 5.3 [1.8, 9.6] 10.8 [6.1, 16.1] 12.4 [8.1, 17.0] 9.3 [5.3, 14.0]
Conjunction 9.1 [4.1, 15.4] 7.6 [4.3, 11.3] 7.7 [4.4, 11.4] 9.4 [5.6, 14.0] 7.5 [3.3, 12.9] 8.0 [3.4, 13.7] 11.7 [7.1, 17.1] 13.2 [8.4, 18.7] 9.3 [5.1, 14.3]
Semantics Concept 5.0 [2.3, 8.4] 8.9 [4.3, 14.2] 5.3 [2.6, 8.7] 6.5 [3.2, 10.5] 4.8 [1.2, 9.3] 3.9 [1.2, 7.5] 8.5 [4.1, 13.7] 9.4 [5.7, 13.7] 6.5 [3.1, 10.7]
Negation 4.8 [2.2, 8.0] 5.2 [2.9, 7.9] 6.5 [3.7, 9.9] 8.2 [4.8, 12.2] 3.9 [1.3, 7.0] 4.6 [2.1, 7.6] 10.5 [6.9, 14.7] 15.2 [10.3, 20.6] 7.4 [4.3, 11.0]
Discourse Disc. markers 4.7 [2.0, 8.0] 1.6 [0.1, 3.6] 5.4 [2.8, 8.3] 2.2 [0.8, 3.9] 1.9 [0.4, 4.0] 3.2 [0.8, 6.1] 6.3 [3.4, 9.5] 8.7 [5.2, 12.9] 4.2 [1.9, 7.0]
Appraisal 5.5 [2.9, 8.6] 2.6 [0.8, 5.0] 4.8 [2.4, 7.7] 3.2 [1.4, 5.3] 3.1 [1.0, 5.8] 2.9 [0.9, 5.5] 7.1 [4.2, 10.4] 14.0 [9.2, 19.2] 5.4 [2.9, 8.4]
Varieties Style 11.8 [6.6, 17.7] 12.2 [7.4, 17.7] 13.1 [8.5, 18.3] 3.6 [1.8, 5.7] 4.8 [2.1, 8.0] 6.4 [2.9, 10.6] 7.4 [4.1, 11.2] 10.9 [6.9, 15.3] 8.8 [5.0, 13.1]
Dialectal 12.6 [8.5, 17.3] 7.4 [4.1, 11.4] 8.4 [5.8, 11.4] 5.9 [3.2, 9.3] 8.9 [4.5, 14.0] 4.7 [2.2, 7.9] 7.0 [4.0, 10.6] 13.2 [8.4, 18.3] 8.5 [5.1, 12.5]
Average 7.4 [4.1, 11.4] 8.6 [5.1, 12.6] 8.3 [5.1, 12.0] 7.5 [4.3, 11.4] 6.7 [3.4, 10.8] 6.4 [3.4, 10.0] 9.3 [5.5, 13.7] 13.0 [8.6, 17.9] 8.4 [4.9, 12.5]
Table 13: NER: Unrobustness (U, %) by model and modification
Category Modification PLM LLM Avg
BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 9.0 [4.8, 16.2] 4.0 [1.6, 9.8] 6.0 [2.8, 12.5] 8.0 [4.1, 15.0] 1.0 [0.2, 5.4] 7.0 [3.4, 13.7] 3.0 [1.0, 8.5] 8.0 [4.1, 15.0] 5.8 [2.8, 12.0]
Geographical 8.0 [4.1, 15.0] 14.0 [8.5, 22.1] 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 4.0 [1.6, 9.8] 5.0 [2.2, 11.2] 6.0 [2.8, 12.5] 10.0 [5.5, 17.4] 8.2 [4.4, 15.2]
Length 19.2 [12.6, 28.0] 15.2 [9.4, 23.5] 14.1 [8.6, 22.3] 15.2 [9.4, 23.5] 20.2 [13.5, 29.2] 17.2 [11.0, 25.8] 8.1 [4.2, 15.1] 13.1 [7.8, 21.2] 15.3 [9.6, 23.6]
Orthography Spelling 11.2 [6.4, 19.0] 3.1 [1.0, 8.6] 7.1 [3.5, 14.0] 5.1 [2.2, 11.4] 2.0 [0.6, 7.1] 4.1 [1.6, 10.0] 7.1 [3.5, 14.0] 4.1 [1.6, 10.0] 5.5 [2.5, 11.8]
Capitalization 15.2 [9.4, 23.5] 14.1 [8.6, 22.3] 7.1 [3.5, 13.9] 7.1 [3.5, 13.9] 4.0 [1.6, 9.9] 5.1 [2.2, 11.3] 2.0 [0.6, 7.1] 6.1 [2.8, 12.6] 7.6 [4.0, 14.3]
Punctuation 1.0 [0.2, 5.5] 7.1 [3.5, 13.9] 1.0 [0.2, 5.5] 1.0 [0.2, 5.5] 3.0 [1.0, 8.5] 1.0 [0.2, 5.5] 2.0 [0.6, 7.1] 4.0 [1.6, 9.9] 2.5 [0.9, 7.7]
Morphology Derivation 4.1 [1.6, 10.0] 3.1 [1.0, 8.6] 5.1 [2.2, 11.4] 4.1 [1.6, 10.0] 1.0 [0.2, 5.6] 1.0 [0.2, 5.6] 5.1 [2.2, 11.4] 2.0 [0.6, 7.1] 3.2 [1.2, 8.7]
Compound 6.2 [2.9, 13.0] 4.2 [1.6, 10.2] 3.1 [1.1, 8.8] 5.2 [2.2, 11.6] 6.2 [2.9, 13.0] 6.2 [2.9, 13.0] 3.1 [1.1, 8.8] 4.2 [1.6, 10.2] 4.8 [2.0, 11.1]
Syntax Voice 35.8 [26.9, 45.8] 34.7 [25.9, 44.7] 41.1 [31.7, 51.1] 26.3 [18.5, 36.0] 15.8 [9.8, 24.4] 30.5 [22.2, 40.4] 9.5 [5.1, 17.0] 15.8 [9.8, 24.4] 26.2 [18.7, 35.5]
Grammar 30.6 [21.1, 42.0] 27.8 [18.8, 39.0] 19.4 [12.0, 30.0] 22.2 [14.2, 33.1] 18.1 [10.9, 28.5] 19.4 [12.0, 30.0] 16.7 [9.8, 26.9] 20.8 [13.1, 31.6] 21.9 [14.0, 32.6]
Conjunction 5.2 [2.2, 11.5] 10.3 [5.7, 17.9] 8.2 [4.2, 15.4] 8.2 [4.2, 15.4] 4.1 [1.6, 10.1] 7.2 [3.5, 14.2] 5.2 [2.2, 11.5] 6.2 [2.9, 12.8] 6.8 [3.3, 13.6]
Semantics Concept 8.0 [4.1, 15.0] 3.0 [1.0, 8.5] 5.0 [2.2, 11.2] 11.0 [6.3, 18.6] 18.0 [11.7, 26.7] 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 10.0 [5.5, 17.4] 9.2 [5.1, 16.4]
Negation 25.5 [17.9, 35.0] 23.5 [16.2, 32.8] 24.5 [17.0, 33.9] 21.4 [14.5, 30.5] 24.5 [17.0, 33.9] 22.4 [15.3, 31.7] 22.4 [15.3, 31.7] 30.6 [22.4, 40.3] 24.4 [17.0, 33.7]
Discourse Disc. markers 8.0 [4.1, 15.0] 6.0 [2.8, 12.5] 8.0 [4.1, 15.0] 19.0 [12.5, 27.8] 10.0 [5.5, 17.4] 7.0 [3.4, 13.7] 9.0 [4.8, 16.2] 11.0 [6.3, 18.6] 9.8 [5.4, 17.0]
Appraisal 5.0 [2.2, 11.2] 8.0 [4.1, 15.0] 4.0 [1.6, 9.8] 8.0 [4.1, 15.0] 7.0 [3.4, 13.7] 9.0 [4.8, 16.2] 5.0 [2.2, 11.2] 7.0 [3.4, 13.7] 6.6 [3.2, 13.2]
Varieties Style 16.0 [10.1, 24.4] 19.0 [12.5, 27.8] 18.0 [11.7, 26.7] 19.0 [12.5, 27.8] 14.0 [8.5, 22.1] 14.0 [8.5, 22.1] 14.0 [8.5, 22.1] 12.0 [7.0, 19.8] 15.8 [9.9, 24.1]
Dialect 8.3 [4.4, 15.1] 26.9 [19.4, 35.9] 22.2 [15.4, 30.9] 24.5 [17.2, 33.7] 11.8 [6.9, 19.4] 16.7 [10.7, 25.1] 2.9 [1.0, 8.3] 12.7 [7.6, 20.6] 15.8 [10.3, 23.6]
Average 12.7 [7.9, 20.3] 13.2 [8.3, 20.8] 11.9 [7.4, 19.3] 12.7 [7.8, 20.4] 9.7 [5.7, 16.8] 10.7 [6.4, 18.0] 7.7 [4.1, 14.5] 10.5 [6.1, 17.8] 11.1 [6.7, 18.5]
Table 14: Coreference: Unrobustness (U, %) by model and modification
A.8 Use Case: Dialogue Contradiction Detection and Sentiment Analysis (extended results)
As Table˜15 shows, on Dialogue contradiction detection task base LLMs are overall more robust to modifications than PLMs; however, reasoning LLMs on average perform similar to LLMs. All models show major brittleness to negation, and the majority of them also lack robustness to change of grammatical roles, which shows that they tend to rely on surface-level, token consistency between the context and the final turn, rather than on logical coherence. Moreover, LLMs (but to a lesser extent PLMs) demonstrate geographical bias, whuile PLMs are particularly brittle to casual style.